OpenStudy (anonymous):

hi please anyone explain me the first question of homework 1

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

did you try to solve it?

OpenStudy (anonymous):

yes but i m not able to solve its calcultion

OpenStudy (anonymous):

The equation tool on this chat box does not allow vector arrows/operators.

OpenStudy (unklerhaukus):

* ``` \[\vec{\mathbf E}(x)=E(x)\hat{\mathbf x} \] ``` \[\vec{\mathbf E}(x)=E(x)\hat{\mathbf x}\]

OpenStudy (unklerhaukus):

\[\vec{\mathbf E}=\vec{\mathbf E}_{Q_1}+\vec{\mathbf E}_{Q_2}\]

OpenStudy (unklerhaukus):

\[=K\frac{Q_1}{r_1^2}\hat{\mathbf x}+K\frac{Q_2}{r_2^2}\hat{\mathbf x}\]\[=K\left(\frac{Q_1}{r_1^2}+\frac{Q_2}{r_2^2}\right)\hat{\mathbf x}\]

OpenStudy (anonymous):

A point charge Q1=−2 μC is located at x=0, and a point charge Q2=+8 μC is placed at x=−0.5 m on the x-axis of a cartesian coordinate system.The goal of this problem is to determine the electric field, E⃗ (x)=E(x)xˆ, at various points along the x-axis. (a)What is E(x) (in N/C) for x=-13.0 m ? i tried which u explain but still my ans wrong :( i dnt know "x" how to deal wd it

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):

\[\hat {\mathbf{r}}_1=-\hat {\mathbf{x}}\]

OpenStudy (unklerhaukus):

we see that \[r_1^2=x^2\] \[r_2^2=(x-0.5)^2\]